OpenStudy (anonymous):

Part 1: Explain, in complete sentences, whether or not the expression x2 – 121 fits one of the special patterns. If so, which one? (2 points) Part 2: Explain how the polynomial is factored. (2 points) Part 3: Provide the factors. (2 points)

OpenStudy (anonymous):

it fits that is (a+b)(a-b)

OpenStudy (anonymous):

\[x^2-121=(x+11)(x-11)\]

OpenStudy (aravindg):

got it @raebaby420 ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

part 1---> it fits in the special series that if 2 numbers like a and b are there so \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (aravindg):

well first of all it is of the form \(a^2-b^2\) so you can immediately split it as \[(a^2-b^2)=(a+b)(a-b)\] this how we factor the polynomial ..with that said can you tell which will be the factors ?